If I use 36 feet of fence and the square footage is 72 square feet what are the lenghts of each side of fence? Thank you
Answer by nerdybill(7384)

If I use 36 feet of fence and the square footage is 72 square feet what are the lenghts of each side of fence?
Let x = width
and y = length
.
then
perimeter = 2(width+length)
36 = 2(x+y)
18 = x+y (equation 1)
.
area = width*length
72 = xy (equation 2)
.
Solve equation 1 for y:
18 = x+y
18-x = y
Substitute above into equation 2 and solve for x:
72 = xy
72 = x(18-x)
72 = 18x-x^2
x^2+72 = 18x
x^2-18x+72 = 0
(x-6)(x-12) = 0
x = {6, 12}
.
Solution:
6 feet by 12 feet